# Many-Shot In-Context Learning in Multimodal Foundation Models

### v1

@@ 4.2 Impact of batching queries
Zero-shot performance improvements from batching queries.  We conduct several additional experiments to investigate why batch querying can lead to large performance improvements under the zero-shot regime on TerraIncognita and UCMerced. We hypothesize that this improvement may be due to three potential benefits from ICL: (1) domain calibration, where the model benefits from seeing more images in the domain in order to adapt to it, (2) class calibration, where seeing images of different classes enables the model to better calibrate its outputs, and (3) self-ICL (shown to be effective in prior work [29]), where the model can learn from self-generated demonstrations due to autoregressive decoding. We design experiments to isolate the potential benefits from each of these types of ICL between asking a single query to batching 50 queries together.
@@ 4.3 Cost and latency analysis
Dataset No Query Batching Query Batching
@@
5 Discussion
In this study, we evaluate many-shot ICL of state-of-the-art multimodal foundation models across 10 datasets and find consistent performance improvements across most of the datasets. Batching queries with many-shot ICL further exhibits substantially reduced per-example latency and inference costs without compromising performance.
Our findings suggest that these multimodal foundation models have the capability of performing ICL with large numbers of demonstrating examples, which may have significant implications on their practical use. For example, it was previously impossible to adapt these large, private models to new tasks and domains, but many-shot ICL would enable users to leverage demonstrating examples to adapt the models. One significant advantage of many-shot ICL is its ability to get quick results even on the same day of model release, and that’s why we can finish our evaluation using GPT-4o within days. Furthermore, fine-tuning open-source models is the standard practice when practitioners have access to moderately sized datasets, but many-shot ICL may remove the need for fine-tuning, making it much easier to develop customized approaches. We note that it remains to be seen how traditional fine-tuning of these models compares to many-shot ICL with foundation models in terms of absolute performance and data efficiency, so future work should explore this. In addition, it is important to study general issues which plague those foundation models, such as hallucinations and biases, under the context of many-shot ICL and batching queries. For example, it would be interesting to explore if carefully curated and large sets of demonstrating examples can reduce biases across different sub-groups. We leave this to future work.
Our study has limitations. First, we only explore performance under many-shot ICL on image classification tasks and with private foundation models. We believe these are the most practically relevant and common multimodal settings, but it is worthwhile for future work to explore potential benefits from many-shot ICL on other tasks and with upcoming open-source multimodal foundation models like LLaMA-3 [30]. Second, even after recent developments to increase context size, the size prohibits many-shot ICL  from being used on datasets with a large number (several hundred or more) of classes. We anticipate that context window sizes will continue to increase in size over time which will mitigate this issue. Third, the datasets which were used to train these private models have not been disclosed, so it is difficult to tell whether the models have been trained on the datasets we selected. We argue that zero-shot performance across the datasets is far from perfect which provides evidence that the datasets have not been used for training, but we cannot determine that with certainty.
@@ 6 Conclusion
In summary, we show that multimodal foundation models are capable of many-shot ICL. We believe that these results pave a promising path forward to improve the adaptability and accessibility of large multimodal foundation models.
@@ Appendix C GPT4(V)-Turbo performance under many-shot ICL
Figure 6: GPT4(V)-Turbo and GPT-4o performance from zero-shot to many-shot ICL. X-axis is in log scale.
@@ D.2 Results
Figure 7: Many-shot ICL performances of medical QA tasks.
Acknowledgments and Disclosure of Funding
We thank Dr. Jeff Dean, Yuhui Zhang, Dr. Mutallip Anwar, Kefan Dong, Rishi Bommasani, Ravi B. Sojitra, Chen Shani and Annie Chen for their feedback on the ideas and manuscript. Yixing Jiang is supported by National Science Scholarship (PhD). This work is also supported by Google cloud credit. Dr. Jonathan Chen has received research funding support in part by NIH/National Institute of Allergy and Infectious Diseases (1R01AI17812101), NIH/National Institute on Drug Abuse Clinical Trials Network (UG1DA015815 - CTN-0136), Gordon and Betty Moore Foundation (Grant #12409), Stanford Artificial Intelligence in Medicine and Imaging - Human-Centered Artificial Intelligence (AIMI-HAI) Partnership Grant, Google, Inc. Research collaboration Co-I to leverage EHR data to predict a range of clinical outcomes, American Heart Association - Strategically Focused Research Network - Diversity in Clinical Trials and NIH-NCATS-CTSA grant (UL1TR003142) for common research resources.

### v2

@@ 4.2 Impact of batching queries
Figure 4: Gemini 1.5 Pro performance under many-shot and zero-shot ICL with varying amount of queries included in every request. We show performance per batch size with the optimal number of demo examples (many-shot) and no demo examples (zero-shot). The $x$-axis is in log scale. Under the many-shot regime, batching queries leads to no substantial drop in performance compared to individual queries when we choose a suitable batch size. For zero-shot, including only one query is suboptimal for many datasets. Error bars are estimated standard deviations using bootstrapping with 1,000 bootstrap replicates.
Zero-shot performance improvements from batching queries.  We conduct several additional experiments to investigate why batch querying can lead to large performance improvements under the zero-shot regime on TerraIncognita and UCMerced. We hypothesize that this improvement may be due to three potential benefits from ICL: (1) domain calibration, where the model benefits from seeing more images in the domain in order to adapt to it, (2) class calibration, where seeing images of different classes enables the model to better calibrate its outputs(Min et al., 2022), and (3) self-ICL (shown to be effective in prior work Chen et al. (2023)), where the model can learn from self-generated demonstrations due to autoregressive decoding. We design experiments to isolate the potential benefits from each of these types of ICL between asking a single query to batching 50 queries together.
@@ 4.3 Cost and latency analysis
5 Discussion
In this study, we evaluate many-shot ICL of state-of-the-art multimodal foundation models across 14 datasets and find consistent performance improvements across most of the datasets. Batching queries with many-shot ICL further exhibits substantially reduced per-example latency and inference costs without compromising performance. We also highlight one future direction for improving open-weights multimodal foundation models as they are unable to learn from demonstrating examples.
Our findings suggest that these multimodal foundation models have the capability of performing ICL with large numbers of demonstrating examples, which may have significant implications on their practical use. For example, it was previously impossible to adapt these large, private models to new tasks and domains, but many-shot ICL would enable users to leverage demonstrating examples to adapt the models. One significant advantage of many-shot ICL is its ability to get quick results even on the same day of model release, and that’s why we can finish our evaluation using GPT-4o and Llama3.2 within days. Furthermore, fine-tuning open-source models is the standard practice when practitioners have access to moderately sized datasets, but many-shot ICL may remove the need for fine-tuning, making it much easier to develop customized approaches. We note that it remains to be seen how traditional fine-tuning of these models compares to many-shot ICL with foundation models in terms of absolute performance and data efficiency, so future work should explore this. In addition, it is important to study general issues which plague those foundation models, such as hallucinations and biases, under the context of many-shot ICL and batching queries. For example, it would be interesting to explore if carefully curated and large sets of demonstrating examples can reduce biases across different sub-groups. We leave this to future work.
Our study has limitations. First, we only explore performance under many-shot ICL on some common multimodal tasks which we believe are the most practically relevant and common multimodal settings, but it is worthwhile for future work to explore potential benefits from many-shot ICL on other tasks such as VL-ICL Bench Zong et al. (2024). Second, even after recent developments to increase context size, the size prohibits many-shot ICL  from being used on datasets with a large number (several hundred or more) of classes. We anticipate that context window sizes will continue to increase in size over time which will mitigate this issue. Third, we only run the ablation experiments on Gemini 1.5 Pro due to budget limit, but it will be interesting to whether these trends hold for GPT-4o. Fourth, the datasets which were used to train these private models have not been disclosed, so it is difficult to tell whether the models have been trained on the datasets we selected. We argue that the zero-shot performance across the datasets is far from perfect, which suggests that the datasets have not been used for training, but we cannot determine that with certainty.
@@ 5 Discussion
Dataset Without Query Batching With Query Batching
@@ 6 Conclusion
In summary, we show that state-of-the-art multimodal foundation models, especially Gemini 1.5 Pro, are capable of many-shot ICLto achieve substantial performance improvements with cost efficiency across multiple domains and tasks. However, open-weights models like Llama 3.2-Vision do not exhibit the same benefits, highlighting a significant future direction. We believe that these results pave a promising path forward to improve the adaptability and accessibility of large multimodal foundation models.
@@ A.3.1 Prefixing images
A.4 Prompt used for visual question answering experiments
prompt = "You’re an expert in answering questions on " + ("radiology"/"satellite")-
+ " images. " + ("Here are some demonstration examples: "-
if num_shot_per_class>0 else "")
for demo in demo_examples:
    prompt += f"""<<IMG>>Given the image above, answer the following question using-
the specified format.
Question: {demo[1]}
Answer: {demo[2]}
    qn_idx = idx + 1
    prompt += f"""<<IMG>>Given the image above, answer the following question using-
the specified format.
Question {qn_idx}: {i.question}
Choices {qn_idx}: {str(i.choices)}
for i in range(start_idx, end_idx):
    qn_idx = i - start_idx + 1
    prompt += f"""
Please respond with the following format for each question:
---BEGIN FORMAT TEMPLATE FOR QUESTION {qn_idx}---
Answer {qn_idx}: [Your Answer Here for Question {qn_idx}]
---END FORMAT TEMPLATE FOR QUESTION {qn_idx}---
A.5 Prompt used for object localization experiments
[…]
for demo in demo_examples:
    prompt += f"""<<IMG>> Return bounding box for the {object_name} in the above
image with this format: [ymin, xmin, ymax, xmax]
[…]
    qn_idx = idx + 1
    prompt += f"""<<IMG>> Return bounding box for the {object_name} in the above
image with this format: [ymin, xmin, ymax, xmax]"""
@@ Appendix C GPT4(V)-Turbo performance under many-shot ICL
Figure 7: GPT4(V)-Turbo and GPT-4o performance from zero-shot to many-shot ICL. We show the accuracy of GPT4(V)-Turbo and GPT-4o as we vary the number of demonstrating examples across the 10 classification datasets. The $x$-axis is in log scale. HAM10000 and TerraIncognita exhibit a relatively smooth log-linear improvement for both models. On the UCMerced, Oxford Pets, and DTD datasets, GPT-4o displays a large drop in performance before reaching the best performance with the largest number of demonstrating examples.
@@ D.2 Results
Figure 8: Many-shot ICL performance of medical QA tasks. The $x$-axis is in log scale. GPT4(V)-Turbo consistently shows better performance compared to Gemini 1.5 Pro. The accuracy tends to increase for GPT4(V)-Turbo, but Gemini 1.5 Pro performance is more variable.